# Diversity of disc viscosities can explain the period ratios of resonant and non-resonant systems of hot super-Earths and mini-Neptunes

### v1

@@ 3.1 Individual system evolution
We show in Fig. 1 the time evolution of a system, in respect to its evolution of semi-major axis, planetary masses, eccentricity and inclination. The planets initially grow by accretion of pebbles, resulting in a smooth increase of the planetary masses. As the planets grow above 1 Earth mass, they start to migrate into the inner disc regions and get caught in resonance configurations. As the pebble isolation mass increases with orbital distance (see appendix E), the planets forming further out in the disc become slightly more massive compared to the inner planets. During their migration, smaller planets can already be scattered.
@@
Once the gas disc dissipates, the planetary eccentricities start to increase to a few percent, while the inclinations increase to around 1 degree for the massive inner planets. The eccentricities of the larger planets remain quite low, resulting in a stable planetary systems for 100 Myr of evolution. Consequently a chain of resonant planets (5:4, 2:1 and 3:2 MMRs) survives in the inner disc.
@@
Fig. 1 shows the evolution of a stable system, which survived until 100 Myr of evolution. Our simulations show two other typical outcomes: Fig. 2 shows a system that underwent instabilities after the gas disc phase that sculpted the inner system, resulting in only 2 remaining inner planets, while Fig. 3 shows the evolution of a system where the inner system is completely removed by scattering caused by the outer planets.
The systems shown in Fig. 2 and Fig. 3 experience strong scattering events after the removal of the gas discs, which leads to large increases of the planetary eccentricities to values around 10% (so around 10 times larger compared to the stable system in Fig. 1).
@@
This is a consequence of the dynamical instabilities in the oute rparts of the system ( AU) that eventually affects the inner system ( AU). Consequently only 2 planets outside of any resonance configuration interior to 1 AU survive (Fig. 2). On the other hand, the instabilities for the system shown in Fig. 3 are so strong that the whole inner system is removed. The cause of the instabilities though, is the same: increase of the orbital eccentricities in the outer system after the gas disc dispersal, leading to an instability that affects and removes the inner system.
3.2 Planetary systems directly after the gas disc-phase
@@
During the first 3 Myr of our simulations, the planets can grow via pebble accretion (and collisions) and migrate, while simultaneously their eccentricity and inclination are damped by gas-disc interactions. Once the gas disc dissipates, only gravitational interactions remain, leading to an increase of eccentricity and inclination with the consequence that instabilities will occur. However, the structure of the systems after the gas disc phase can already give first hints into the differences of low vs. high viscosity planet formation environments.
@@
In Fig. 4 we show the number of planets that have formed at the end of the gas disc phase (left), as well as their period ratios (right). We also show the systems that could be identified by observations as well as the synthetic observations of the simulations of Izidoro et al. (2021). In addition we also show the observed Kepler sample (thin grey lines) as well as the planetary chains observed by Kepler (thick grey lines). We note that we only consider planets interior of 0.7 AU to be observable, even though more planets can be hidden exterior to that (see appendix C). In addition, we only include planets above 1 Earth mass in this statistics, because smaller planets are very likely to be missed in transit observations (e.g. Mulders et al. 2018)
@@
Within 0.7 AU, our low viscosity simulations show a maximum of 6 planets, while the high viscosity simulations of Izidoro et al. (2021) show systems with up to 10 planets. This difference can be explained by the slightly different set-ups used here. In particular, Izidoro et al. (2021) used 60 planetary embryos per simulations, allowing more planets to be present in the first place. Another source of difference is the migration velocity of the growing planets, determined by the disc’s viscosity. Discs with lower viscosity allow early (partial) gap opening, which results in a transition of the planets into a slower migration regime compared to their high viscosity counterparts (see appendix A). Consequently not as many planets from the outer disc migrate interior to 0.7 AU, resulting in different systems.
@@
While the systems in our simulations have always at least 2 planets with 1 Earth mass or more, observationally, the picture is different. Due to the mutual inclinations between the planets, nearly 30% of all systems would only show one transiting planets, while the number of systems with 5 transiting planets is significantly reduced compared to the pure simulation data. This is in line with the results of Izidoro et al. (2021), who have observed the same behaviour in their simulations. Comparing to the observational constraints from Kepler clearly shows that the number of single transiting planets is not reproduced at the stage directly after gas disc dissipation, but requires some instabilities to increase the levels of mutual inclinations and to reduce the number of planets (see below).
@@
The period ratios on the other hand show that most of the formed planets are in resonant configurations at the end of the gas disc lifetime, cause by inward migration and trapping at the disc’s inner edge. This is, of course, in contrast to the observed period ratios by Kepler, where only a small fraction of systems is supposed to be in a resonant configuration. This result is in line with Izidoro et al. (2021), where instabilities are necessary to break the resonant configuration, as we also study (see below). However, there is a significant difference in the observed distribution of the period ratios between our here presented simulations and the ones from Izidoro et al. (2021): the resonant chains in Izidoro et al. (2021) are much more compact, compared to our here presented simulations.
@@
The difference in the period ratios of the systems can again be explained by planetary migration. As the planets migrate faster in the Izidoro et al. (2021) simulations, they can naturally come closer to each other compared to the simulations in a low viscosity environment. In particular, the simulations by Izidoro et al. (2021) seem to overestimate planets in the 7:6 and 5:4 MMRs, while simultaneously underestimating the planets in the 3:2 and 2:1 MMRs. In contrast, the low viscosity simulations show a much better match at these MMRs. We show in appendix D that this difference is truly related to the disc’s viscosity rather than the number of embryos. On the other hand, we discussed here just the simulations at the end of the gas-disc lifetime, and instabilities are expected to change these configurations. We will discuss this next.
@@ 3.3 Planetary systems after the instability phase
After the gas-phase, the planets can only interact gravitationally and lose the damping effects of the disc on eccentricity and inclination. Consequently, the planetary systems become unstable and change their orbital configurations. In the following, we classify as chains, systems that have not undergone a major re-arrangement after the gas-disc phase.
@@
In Fig. 5 we show the number of detected planets (left) and their period ratios (right) after 100 Myr of system evolution from our simulations. As before we also show the distribution for the detected systems as well as the results from Izidoro et al. (2021) in comparison to the Kepler data.
@@
As a consequence of the instability, many planetary systems have no inner system left and therefore show no inner planets (see appendix C). Consequently, also the number of systems with 3 or more inner planets are greatly reduced. However, the reduced number and increased mutual inclinations of systems allow a much better match to the Kepler observations in respect to systems with only one transiting planet, in line with the results of Izidoro et al. (2021). Also the systems with multiple transiting planets match much better after the instability phase.
@@
The resulting period ratios are determined by the systems with at least 2 surviving planets (limiting our statistics to some level, see appendix C). The general results of our low viscosity simulations do not match the period distribution from the Kepler systems, as the systems in our simulations remain too tightly packed compared to the Kepler systems, even when applying synthetic observations. In contrast, the simulations of Izidoro et al. (2021) give a much better match to the period ratio distribution.
@@
The differences in these results could be explained by different strength of instabilities between the different system configurations. While the instabilities in our simulations mostly result in a removal of planets from the system, the instabilities in the simulations of Izidoro et al. (2021) mostly show a re-arrangement of the system with more survivors giving a better period-ratio distribution compared to the observations. This is caused by the difference in the outer system architecture. While the systems formed in the high viscosity environments show basically no massive planets exterior to 200 days due to the fast inward migration of planets of a few Earth masses, the opposite is the case at low viscosity. In this case, even the more massive planets can not migrate to the inner disc (see appendix A), resulting in outer systems with several planets above a few Earth masses. In combination with the lower planetary masses of the inner systems in the low viscosity case (see appendix E), the instabilities triggered by the outer systems results more likely in the removal of the inner planets rather than in just a dynamical re-arrangement. Using a mixture of 2% stable and 98% unstable systems (as done in Izidoro et al. (2021) to achieve the best match to observations) does not allow for a good match to the overall period ratios (see appendix F) of the low viscosity simulations as well.
@@
In Fig. 6 we plot the number of observed planets and their period distributions only for the chains formed in our simulations, as well as the chains from Izidoro et al. (2021) after 100 Myr of evolution. The chains from our simulations harbour only 4 or more planets in contrast to the Kepler observations and the simulations of Izidoro et al. (2021), who also show chains with a smaller number of planets. In addition the high viscosity simulations predict chains with 7 or more planets, which have not been detected in such abundance, yet, but should exist (e.g. Trappist-1).
@@
On the other hand, the surviving chains show a wider period ratio distribution for the low viscosity simulations compared to the high viscosity simulations. In particular the low viscosity simulations show more planets in the 3:2 and 2:1 configurations compared to the high viscosity simulations, giving a better agreement to the observational constraints from the chains alone.
@@
The eccentricities of the planets are shaped by the gravitational interactions between the bodies, where more violent instabilities can cause larger eccentricities. Consequently, the eccentricity distributions of the two sets of simulations (Fig. 7) are different. The low viscosity simulations experience more violent instabilities, caused by the outer planets, while the higher viscosity simulations do not feature outer planets sufficiently close to the inner system and sufficiently massive enough to affect it. Consequently, they experience only “mild” instabilities, resulting in the fact that the eccentricities of the planets in the low viscosity simulations are larger.
@@
In contrast, the eccentricity distribution of the surviving planets in chains (right in Fig. 7) is very similar. This is caused by the fact that chains, if they want to survive, should also somehow avoid being strongly affect by external planets and thus keep their eccentricities low. In the planetary systems shown in Fig. 3 and Fig. 2, the eccentricities increase above 10%, resulting in violent instabilities that break the chains. In contrast, systems that remain in a resonant configuration have low eccentricities (Fig. 1), as the low eccentricities prevent orbit crossing and instabilities in the first place.
@@
We note that, in principle, more compact resonance configurations (e.g. 7:6, 6:5, 5:4) should result in more violent instabilities compared to wider resonance configurations (e.g. 3:2, 2:1). An example for a well described less violent instability would be the dynamical instability of the solar system in the Nice model, even though it features more massive planets on much larger orbits (e.g. Tsiganis et al. 2005; Nesvorný 2018). In our case, instabilties tend to be more violent not necessarily due to the compactness of the resonant configurations themselves, but instead due to the effect of external perturbers (see Fig. 3 and Fig. 2). We will explore this mechanism in more detail in future work.
@@ 4 Discussion
We showed that the outcome of our simulations depends crucially on the disc’s viscosity (see appendix D). We now discuss the influence of the different (changed) parameters on the outcome of our simulations.
@@ 4.1 Planetary migration
The planetary migration rates are not only set by the disc’s viscosity (e.g. Paardekooper et al. 2011), but also by the disc’s properties, especially the surface density and the temperature (which in turn set the entropy). We used here the migration formulae by Paardekooper et al. (2011), which allows reduced (or even outward) migration rates at higher viscosities, depending on the disc’s profile. We have used here the disc profile of Bitsch et al. (2015a), where the disc has already evolved for 2 Myr. At this stage, the inner disc’s aspect ratio is rather small (H/r 0.025-0.03), resulting in an easier partial gap opening of small planets (Crida & Morbidelli, 2007; Bitsch et al., 2013; Kanagawa et al., 2018; Pichierri et al., 2023), which reduces their inward migration speed compared to simulations at higher viscosity (see appendix A), where gap opening is hindered. Consequently the exact outcomes of our simulations are influenced by the disc profile that we have chosen.
We also note that other migration formulae (e.g. Jiménez & Masset 2017) give different migration rates, where especially the heating torque could change the migration direction of fast growing small mass planets (e.g. Benítez-Llambay et al. 2015; Masset 2017). However, the heating torque diminishes once the planets reach the pebble isolation mass, resulting nominal inward migration at low viscosity. The overall outcome of planet formation simulations is not influenced greatly by the heating torque or different migration prescriptions (Baumann & Bitsch, 2020), however the detailed migration rates and disc parameters matter for the formation of specific systems (e.g. Hühn et al. 2021).
The relative migration velocities of the planets set in which resonance planets can be trapped (e.g. Huang & Ormel 2023). This is independent of the cause of the migration rates, be it disc profile or viscosity. We here explored how low viscosities influence the migration rates. A similar effect could also be achieved via different disc parameters like a lower overall disc surface density. However, this would also influence the planetary growth rates via pebble accretion as well as the damping rates of eccentricity and inclination, making direct comparisons and drawing definite conclusions more difficult.
In order to test the influence of the disc’s viscosity on resonance trapping, we show in appendix G the outcome of simulations of two migrating equal mass planets in different disc viscosity environments. While the exact outcomes of the simulations depends on the disc parameters, it is clear that the viscosity - which sets the migration rate - determines the resonance that the planets are trapped in. Planets migrating in high viscosity environments are trapped in closer resonances compared to planets migrating in low viscosity environments.

### v2

@@ 3.1 Individual system evolution
Figure 1 shows the time evolution of a system with respect to its evolution in semi-major axis, planetary masses, eccentricity, and inclination. The planets initially grow by accretion of pebbles, resulting in a smooth increase in the planetary masses. As the planets grow above 1 Earth mass, they start to migrate into the inner disc regions and get caught in resonance configurations. As the pebble-isolation mass increases with orbital distance (see Appendix E), the planets forming further out in the disc become slightly more massive compared to the inner planets. During their migration, smaller planets can already be scattered.
@@
Once the gas disc dissipates, the planetary eccentricities start to increase to a few percent, while the inclinations increase to around 1 degree for the massive inner planets. The eccentricities of the larger planets remain quite low, resulting in stable planetary systems for 100 Myr of evolution. Consequently a chain of resonant planets (5:4, 2:1 and 3:2 MMRs) survives in the inner disc.
@@
Figure 1 shows the evolution of a stable system, which survived until 100 Myr of evolution. Our simulations show two other typical outcomes: Fig. 2 shows a system that underwent instabilities after the gas-disc phase that sculpted the inner system, resulting in only two remaining inner planets, while Fig. 3 shows the evolution of a system where the inner system is completely removed by scattering caused by the outer planets.
@@
The systems shown in Figs. 2 and 3 experience strong scattering events after the removal of the gas discs, which leads to large increases in the planetary eccentricities to values of around 10% (i.e. around ten times larger than the stable system in Fig. 1). This is a consequence of the dynamical instabilities in the outer parts of the system ( AU), which eventually affect the inner system ( AU). Consequently, only two planets survive outside of any resonance configuration interior to 1 AU (Fig. 2). On the other hand, the instabilities for the system shown in Fig. 3 are so strong that the whole inner system is removed. However, the cause of the instabilities is the same: an increase in the orbital eccentricities in the outer system after the gas disc dispersal, leading to an instability that affects and removes the inner system.
3.2 Planetary systems directly after the gas disc phase
@@
During the first 3 Myr of our simulations, the planets can grow via pebble accretion (and collisions) and migrate, while simultaneously their eccentricity and inclination are damped by gas–disc interactions. Once the gas disc dissipates, only gravitational interactions remain, leading to an increase in eccentricity and inclination with the consequence that instabilities will occur. However, the structure of the systems after the gas-disc phase can already provide some initial hints as to the differences between low- and high-viscosity planet formation environments.
@@
Fig. 4 shows the number of planets that have formed at the end of the gas-disc phase (left), as well as their period ratios (right). We also show the systems that could be identified by observations as well as the synthetic observations of the simulations of Izidoro et al. (2021). In addition, we show the observed Kepler sample (thin grey lines) as well as the planetary chains observed by Kepler (thick grey lines). We note that we only consider planets interior to 0.7 AU as observable, even though more planets can be hidden exterior to that (see Appendix C). In addition, we only include planets above 1 Earth mass in our statistical analysis, because smaller planets are very likely to be missed in transit observations (e.g. Mulders et al. 2018)
@@
Within 0.7 AU, our low-viscosity simulations show a maximum of six planets, while the high-viscosity simulations of Izidoro et al. (2021) show systems with up to ten planets. This difference can be explained by the slightly different setups used here. In particular, Izidoro et al. (2021) used 60 planetary embryos per simulation, allowing more planets to be present in the first place. Another difference is the migration velocity of the growing planets, which is determined by the disc viscosity. Discs with lower viscosity allow early (partial) gap opening, which results in a transition of the planets into a slower migration regime compared to their high-viscosity counterparts (see Appendix A). Consequently, not as many planets from the outer disc migrate interior to 0.7 AU, resulting in different systems.
@@
While the systems in our simulations always have at least two planets with 1 Earth mass or more, observationally, the picture is different. Due to the mutual inclinations between the planets, nearly 30% of all systems would only show one transiting planet, while the number of systems with five transiting planets is significantly reduced compared to the pure simulation data. This is in line with the results of Izidoro et al. (2021), who observed the same behaviour in their simulations. Comparing to the observational constraints from Kepler clearly shows that the number of single transiting planets is not reproduced at the stage directly after dissipation of the gas disc; this requires some instability to increase the level of mutual inclination and to reduce the number of planets (see below).
@@
The period ratios on the other hand show that most of the formed planets are in resonant configurations at the end of the gas-disc lifetime, which is caused by inward migration and trapping at the inner edge of the disc. This is, of course, in contrast to the observed period ratios by Kepler, where only a small fraction of systems are expected to be in a resonant configuration. This result is in line with the findings of Izidoro et al. (2021), where instabilities are necessary to break the resonant configuration; we study this aspect further below. However, there is a significant difference in the observed distribution of the period ratios between the simulations presented here and those of Izidoro et al. (2021): the resonant chains in Izidoro et al. (2021) are much more compact than those in the simulations presented here.
@@
The difference in the period ratios of the systems can again be explained by planetary migration. As the planets migrate faster in the Izidoro et al. (2021) simulations, they can naturally come closer to each other than they are able to in simulations in a low-viscosity environment. In particular, the simulations by Izidoro et al. (2021) seem to overestimate planetary systems in the 7:6 and 5:4 MMRs, while simultaneously underestimating the planets in the 3:2 and 2:1 MMRs. In contrast, the low-viscosity simulations show a much better match at these MMRs. We show in Appendix D that this difference is truly related to disc viscosity rather than the number of embryos. On the other hand, here we only discuss the simulations at the end of the gas-disc lifetime, and instabilities are expected to change these configurations. We discuss this aspect further in the following section.
@@ 3.3 Planetary systems after the instability phase
After the gas phase, the planets can only interact gravitationally and are no longer influenced by the damping effects of the disc that previously acted on their eccentricity and inclination. Consequently, the planetary systems become unstable and their orbital configurations change. In the following, we classify systems that have not undergone a major rearrangement after the gas-disc phase as ‘chains’.
@@
Fig. 5 shows the number of detected planets (left) and their period ratios (right) in our simulations after 100 Myr of system evolution. As before, we also show the distribution of the detected systems as well as the results from Izidoro et al. (2021) in comparison to the Kepler data.
@@
As a consequence of the instability, many planetary systems have no inner system left and therefore show no inner planets (see Appendix C). Consequently, the number of systems with three or more inner planets is also greatly reduced. However, the reduced number and increased mutual inclination of the systems allow a much better match to the Kepler observations regarding the systems with only one transiting planet, in line with the results of Izidoro et al. (2021). The systems with multiple transiting planets also match much more closely after the instability phase.
@@
The resulting period ratios are determined by the systems with at least two surviving planets (limiting our statistics to some level; see Appendix C). The general results of our low-viscosity simulations do not match the period distribution from the Kepler systems, as the systems in our simulations remain too tightly packed compared to the Kepler systems, even when applying synthetic observations. In contrast, the simulations of Izidoro et al. (2021) are a much better match to the Kepler systems in terms of their period ratio distribution.
@@
The differences in these results could be explained by the different strengths of the instabilities between the different system configurations. While the instabilities in our simulations mostly result in a removal of planets from the system, the instabilities in the simulations of Izidoro et al. (2021) mostly show a rearrangement of the system, with more survivors, which leads to period ratio distributions that are in better agreement with observations. This is caused by the difference in the outer system architecture. While the systems formed in the high-viscosity environments show almost no massive planets exterior to 200 days due to the fast inward migration of planets of a few Earth masses, the opposite is true at low viscosity, where even the more massive planets cannot migrate to the inner disc (see Appendix A), resulting in outer systems with several planets above a few Earth masses. In combination with the lower planetary masses of the inner systems in the low-viscosity case (see Appendix E), the instabilities triggered by the outer systems are more likely to result in the removal of the inner planets rather than simply a dynamical rearrangement. Using a mixture of 2% stable and 98% unstable systems (as done in Izidoro et al. (2021) to achieve the best match to observations) does not allow a good match to the overall period ratios (see Appendix F) of the low-viscosity simulations as well.
@@
In Fig. 6 we plot the number of observed planets and their period distributions only for the chains formed in our simulations, as well as the chains from Izidoro et al. (2021) after 100 Myr of evolution. The chains from our simulations harbour only four or more planets, in contrast to the Kepler observations; this is also different from the simulations of Izidoro et al. (2021), who also show chains with a smaller number of planets. In addition, the high-viscosity simulations predict chains with seven or more planets, which have not yet been detected in such abundance, but should exist in theory (e.g. Trappist-1).
@@
On the other hand, the surviving chains show a wider period ratio distribution in the low-viscosity simulations than in the high-viscosity simulations. In particular, the low-viscosity simulations show more planets in the 3:2 and 2:1 configurations compared to the high-viscosity simulations, which leads to better agreement with the observational constraints from the chains alone.
@@
The eccentricities of the planets are shaped by the gravitational interactions between the bodies, where more violent instabilities can cause larger eccentricities. Consequently, the eccentricity distributions of the two sets of simulations (Fig. 7) are different. The low-viscosity simulations experience more violent instabilities, which are caused by the outer planets, while the higher viscosity simulations do not feature outer planets that are both sufficiently massive and sufficiently close to the inner system to affect it. Consequently, they experience only ‘mild’ instabilities, resulting in planets with larger eccentricities in the low-viscosity simulations.
@@
In contrast, the eccentricity distribution of the surviving planets in chains (right in Fig. 7) is very similar. This is caused by the fact that chains, if they are to survive, must also somehow avoid being strongly affected by external planets and must therefore keep their eccentricities low. In the planetary systems shown in Fig. 3 and Fig. 2, the eccentricities increase above 10%, resulting in violent instabilities that break the chains. In contrast, systems that remain in a resonant configuration have low eccentricities (Fig. 1), as the low eccentricities prevent orbit crossing and instabilities in general.
@@
We note that, in principle, more compact resonance configurations (e.g. 7:6, 6:5, 5:4) should result in more violent instabilities compared to wider resonance configurations (e.g. 3:2, 2:1). An example of a well-described less violent instability would be the dynamical instability of the Solar System in the Nice model, although it features more massive planets on much larger orbits (e.g. Tsiganis et al. 2005; Nesvorný 2018). In the cases analysed here, instabilities tend to be more violent, not necessarily due to the compactness of the resonant configurations themselves, but instead due to the effect of external perturbers (see Fig. 3 and Fig. 2). We will explore this mechanism in more detail in future work.
@@ 4 Discussion
We show that the outcome of our simulations depends crucially on the viscosity of the disc (see Appendix D). We now discuss the influence of the different (changed) parameters on the outcomes of our simulations.
@@ 4.1 Planetary migration
The planetary migration rates are not only set by the viscosity of the disc (e.g. Paardekooper et al. 2011), but also by its other properties, and in particular its surface density and temperature (which in turn set the entropy). Here we use the migration formulae by Paardekooper et al. (2011), which allow reduced (or even outward) migration rates at higher viscosities, depending on the disc profile. We also use the disc profile of Bitsch et al. (2015a), where the disc has already evolved for 2 Myr. At this stage, the aspect ratio of the inner disc is relatively small (H/r 0.025-0.03), resulting in an easier partial gap opening by small planets (Crida & Morbidelli, 2007; Bitsch et al., 2013; Kanagawa et al., 2018; Pichierri et al., 2023), which reduces their inward migration speed compared to simulations at higher viscosity (see Appendix A), where gap opening is hindered. Consequently, the exact outcomes of our simulations are influenced by our choice of disc profile.
We also note that other migration formulae (e.g. Jiménez & Masset 2017) give different migration rates, where the heating torque in particular can change the migration direction of fast-growing small-mass planets (e.g. Benítez-Llambay et al. 2015; Masset 2017). However, the heating torque diminishes once the planets reach the pebble-isolation mass, resulting in nominal inward migration at low viscosity. The overall outcome of planet-formation simulations is not greatly influenced by the heating torque or different migration prescriptions (Baumann & Bitsch, 2020); however, the detailed migration rates and disc parameters are important in the formation of specific systems (e.g. Hühn et al. 2021).
The relative migration velocities of the planets set the resonances in which planets can be trapped (e.g. Huang & Ormel 2023). This is independent of the factor that sets the migration rates, be it disc profile or viscosity. Here we explore how low viscosities influence the migration rates. A similar effect could also be achieved via different disc parameters, such as a lower overall disc surface density. However, this would also influence the rate at which planets grow via pebble accretion, as well as the damping rates of eccentricity and inclination, making it more difficult to make direct comparisons and draw robust conclusions.
In order to test the influence of disc viscosity on resonance trapping, we show in Appendix G the outcomes of simulations of two migrating equal-mass planets in different disc-viscosity environments. While the exact outcomes of the simulations depend on the disc parameters, it is clear that the viscosity —which sets the migration rate— determines the resonance that the planets are trapped in. Planets migrating in high-viscosity environments are trapped in closer resonances compared to planets migrating in low-viscosity environments.